# Membrane phononic integrated circuits

### v1

@@ XII Conclusion
In the future we anticipate far more circuit designs could be imagined and realised using the theory presented in this tutorial. For instance, one direction of future work may be creating networks of coupled resonators, using evanescent couplers to achieve precisely tailored coupling strengths, and electrostatic actuation or geometric elongation to introduce Duffing nonlinearity. The ability to fabricate nearly arbitrary network topologies with precisely tunable coupling strengths suggests membrane phononics may be a potent platform for experimenting with systems of nonlinear oscillators. Example applications include mechanical error correction [21, 62], machine learning and neuromorphic computing [155, 156] and basic research into complex networks [157, 158, 159, 158].
Appendix A Parameters and material properties
@@
Volumetric density, $\rho$ $3200\,\mathrm{kg\,m^{-3}}$ [160]
@@
Table 2: Typical material properties corresponding to high tensile-stress thin-film silicon nitride

### v2

@@ XII Conclusion
In the future we anticipate far more circuit designs could be imagined and realised using the theory presented in this tutorial. For instance, one direction of future work may be creating networks of coupled resonators, using evanescent couplers to achieve precisely tailored coupling strengths, and electrostatic actuation or geometric elongation to introduce Duffing nonlinearity. The ability to fabricate nearly arbitrary network topologies with precisely tunable coupling strengths suggests membrane phononics may be a potent platform for experimenting with systems of nonlinear oscillators. Example applications include mechanical error correction [21, 69], machine learning and neuromorphic computing [166, 167, 168] and basic research into complex networks [169, 170, 171, 170].
Acknowledgements
This research was primarily funded by the Australian Research Council and Lockheed Martin Corporation through the Australian Research Council Linkage Grant No. LP190101159. Support was also provided by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (No. CE170100009). G.I.H. (No. DE210100848) and C.G.B. (No. FT240100405) acknowledge their Australian Research Council Fellowships.
The authors would like to thank Rachpon Kalra and George Brawley for helpful discussions in the development of this work.
Appendix A Appendix: Parameters and material properties
Table 2 lists material properties that we use to describe high tensile-stress thin-film silicon nitride membranes.
@@ Appendix A Appendix: Parameters and material properties
Volumetric density, $\rho$ $3200\,\mathrm{kg\cdot m^{-3}}$ [172]
@@
Table 2: Typical material properties used in simulations, corresponding to high tensile-stress thin-film silicon nitride.